3a+4b=9
-3a-2b=-3

Answer by ReadingBoosters(3246)

When the variable is the same in two equation but one is negative, add them together.
3a + -3a + 4b + -2b = 9 + -3 ==> the a variable will cancel, solve for b
2b= 6 ==> divide both sides by 2
b = 3
...
Substitute the value of b into either equation and solve for a
3a + 4(3) = 9
3a + 12 = 9 ==> isolate the a variable
3a = 9 - 12
3a = -3 ==> divide both sides by 3
a = -1
...
Solutions
a=-1, b=3
..............................
Have more questions?
HomeworkHelpers@ReadingBoosters.com.
Delighted to help!
-Reading Boosters
Website: www.MyHomeworkAnswers.com